OpenStudy (anonymous):

Factor completely. 4y^2-36y+72

jimthompson5910 (jim_thompson5910):

hint: 4y^2-36y+72 = 4(y^2 - 9y + 8)

jimthompson5910 (jim_thompson5910):

oops 4y^2-36y+72 = 4(y^2 - 9y + 18)

OpenStudy (nincompoop):

first see if everything is divisible by 4 then try the rest

OpenStudy (anonymous):

Would it be something like: (4y-6+?)(4y-6y+?)

OpenStudy (anonymous):

Everything is divisible by 4!

jimthompson5910 (jim_thompson5910):

try to factor y^2 - 9y + 18

jimthompson5910 (jim_thompson5910):

when you factor that 4 out, you can temporarily ignore it

OpenStudy (anonymous):

is it (y^2-9y+18) ?? or would i have to divide that into two?

jimthompson5910 (jim_thompson5910):

find two numbers that multiply to 18 and add to -9

OpenStudy (anonymous):

6 and 3

jimthompson5910 (jim_thompson5910):

close

jimthompson5910 (jim_thompson5910):

-6 and -3

OpenStudy (anonymous):

-6 and -3

jimthompson5910 (jim_thompson5910):

-6 plus -3 = -9 -6 times -3 = 18

jimthompson5910 (jim_thompson5910):

it's always good to check

jimthompson5910 (jim_thompson5910):

so that means y^2 - 9y + 18 factors to (y - 6)(y - 3) notice how we have the -6 and -3 in the factorization

OpenStudy (anonymous):

Thank you for the help. It so much clearer now!

jimthompson5910 (jim_thompson5910):

so you should have 4(y - 6)(y - 3) as your final answer don't forget that 4 we pulled out for a short while

jimthompson5910 (jim_thompson5910):

you're welcome